Determine the number of moles of the compound and determine the number of moles of each type of atom in each of the following.

(a) 2.28 g of potassium bromide, KBr

compound_______mol

K______________mol

Br_____________mol

(b) 0.1429 g of hydrogen(I) phosphite, H3PO3

compound____________mol

H__________________mol

P_________________mol

O________________mol

(c) 26.0 kg of calcium carbonate, CaCO3

compound__________mol

Ca________________mol

C_________________mol

O_________________mol

(d) 78.36 g of aluminum sulfate, Al2(SO4)3

compound_________mol

Al________________mol

S________________mol

O_______________mol

(e) 0.1287 mg of xanthine, C5H4N4O2

compound____________mol

C___________________mol

H__________________mol

N_________________mol

O_________________mol
Answer #1

a. no of moles of KBr     = W/G.M.Wt

                                         = 2.28/119 = 0.0192 moles

compound   0.0192moles

K               0.0192 moles

Br            0.0192moles

b. no of moles of H3PO3     = W/G.M.Wt

                                           = 0.1429/82 = 0.00174moles

compound      0.00174 moles

H                      0.00522 moles

P                       0.00174 moles

O                     0.00522 moles

c. no of moles of CaCO3     = W/G.M.Wt

                                            = 26000/100 = 260 moles

compound     260 moles

Ca                 260 moles

C                  260 moles

O                  780 moles

d. no of moles of Al2(SO4)3    = W/G.M.Wt

                                                    = 78.36/342 = 0.229 moles

compound     0.229 moles

Al                  0.458 moles

S                   0.687 moles

O                 2.748 moles

e. no of moles of C5H4N4O2 = W/G.M.Wt

                                                 = 0.1287*10^-3/152   = 8.467*10^-7 moles

compound       8.467*10^-7 moles

C                       4.2335*10^-6 moles

H                        3.3868*10^-6 moles

N                       3.3868*10^-6 moles

O                       1.6934*10^-6 moles